11. For arbitrary constants c1,c2, And c3 show that an=c12n+c25n+c3n5n satisfies the recurrence relations an-12an-1+45an-2-50an-3=0


1
Expert's answer
2022-03-05T08:07:48-0500

To prove "a_n=c_12^n+c_25^n+c_3n5^n"  satisfies the recurrence relation "a_n-12a_{n-1}+45a_{n-2}-50a_{n-3}=0", it is enough to show that the solution of the recurrence relation is "a_n=c_12^n+c_25^n+c_3n5^n".


The auxiliary equation of the recurrence relation is, "r^3-12r^2+45r - 50 =0."

Using synthetic division,


"\begin{array}{l|lll}
2&1&-12&~~~45&-50\\
& 0 & ~~~~~2 & -20 & ~~~50\\
\hline
& 1 & -10 & ~~~25 & \mid 0\\
\hline
\end{array}"


We get 2 is a root of the equation. Hence,

"\begin{aligned}
0 = r^3-12r^2+45r - 50 &= (r-2)(r^2-10r+25)\\
&=(r-2)(r-5)^2
\end{aligned}"

Thus the roots of the auxiliary equation are, "r=2,5,5".

The solution of the recurrence relation is

"a_{n} = c_12^n + (c_2 + c_3n)5^{n} = c_12^n + c_25^{n} + c_3n5^{n}". Hence "a_{n}" satisfies the given recurrence relation.